# Proving the Limited Scalability of Centralized Distributed Optimization via a New Lower Bound Construction

### v1

@@ B.3 Proof of lemmas
Hence, the worker receives a stream of coordinates $\nu_{1},\nu_{2},\dots$ such that $\mathbb{P}\left(\nu_{i}\in[K]\middle|\nu_{1},\dots,\nu_{i-1}\right)\leq\frac{K}{d-i+1}$ for all $i\leq d.$ Let $\mu_{1,i,1}$ be the number of received coordinates until the moment when the last received coordinate belongs to $[K]$ in worker $i.$ Similarly, let $\mu_{1,i,k}$ be the number of received coordinates until the moment when the last received coordinate belongs to $[K]$, after the $(k-1)$th time this has happened in worker $i.$ In other words, worker $i$ should receive $\mu_{1,i,1}$ coordinates to obtain a coordinate that belongs to $[K]$. To get the next coordinate that belongs to $[K]$, the worker should receive $\mu_{1,i,2}$ coordinates, and so on. Then,
@@
$t_{1}\geq\min_{i\in[n]}\left\{\min\left\{h\sum_{k=1}^{\frac{K}{2}}\eta_{1,i,k},\tau_{\textnormal{s}}\sum_{k=1}^{\frac{K}{2}}\mu_{1,i,k}\right\}\right\},$
@@
More formally, $\eta_{2,i,k}$ can be defined as the number of stochastic gradient computations until the first moment when a coordinate is not zeroed out in (26), after the moment when the $(k-1)$th coordinate is no longer zeroed out, when $\textnormal{prog}^{K}$ of the input points to the stochastic gradients is $\geq K,$ and $\mu_{2,i,k}$ be the number of received coordinates until the moment when the last received coordinate belongs to $\{K+1,\dots,2K\}$, after the $(k-1)$th time this has happened, when $\textnormal{prog}^{1}$ of the input points to the compressor is $\geq K+1,$ and so on.
@@
C.1 Main Concentration Lemma
See C.1
@@ C.1 Main Concentration Lemma
where we bound $\max$ by $\sum$.∑ . Using the temporal definitions of $\{\beta_{B,i}\},$
@@
Substituting the last inequality to (36) and repeating the steps $B-1$ more times, we get
@@
Appendix D Main Theorem with Worker-to-Server Communication
Protocol 2
1:  Input: Algorithm $A\in\mathcal{A}_{\textnormal{zr}}$
2:  Init $I=\emptyset$ (all available information) on the server
5:  for $i=1,\dots,n$ (in parallel on the workers) do
6:     for $k=0,1,\dots$ do
7:        Algorithm $A$ calculates a new point $x$ based on local information $I_{i}$ (takes 00 seconds)
8:        Calculate a stochastic gradient $\nabla f(x;\xi),$ $\quad\xi\sim\mathcal{D}_{\xi}$ (takes $h$ seconds)
9:        Add $\nabla f(x;\xi)$ to $I_{i}$ (takes 00 seconds)
10:     end for
11:  end for
12:  for $i=1,\dots,n$ (in parallel on the workers) do
13:     for $k=0,1,\dots$ do
14:         Algorithm $A$ calculates new points $\{\bar{s}^{k}_{i}\}$ based on local information $I_{i}$ (takes 00 seconds)
15:         Send $\bar{\mathcal{C}}^{k}_{i}(\bar{s}^{k}_{i})$ to the server: add to $\bar{\mathcal{C}}^{k}_{i}(\bar{s}^{k}_{i})$ to $I$ (takes $\tau_{\textnormal{w}}\times\bar{P}^{k}_{i}$ seconds, where $\bar{P}^{k}_{i}$ is the number of coordinates retained by $\bar{\mathcal{C}}^{k}_{i}(\bar{s}^{k}_{i})$)
16:     end for
17:  end for
18:  for $i=1,\dots,n$ (in parallel on the server) do
19:     for $k=0,1,\dots$ do
20:        Algorithm $A$ calculates new points $s_{i}^{k}$ based on local information $I$ (takes 00 seconds)
21:        Algorithm $A$ compresses the points $\mathcal{C}^{k}_{i}(s_{i}^{k})$ (takes 00 seconds)
22:        Send $\mathcal{C}^{k}_{i}(s_{i}^{k})$ to the worker $i$: add to $\mathcal{C}^{k}_{i}(s_{i}^{k})$ to $I_{i}$ (takes $\tau_{\textnormal{s}}\times P^{k}_{i}$ seconds, where $P^{k}_{i}$ is the number of coordinates retained by $\mathcal{C}^{k}_{i}(s_{i}^{k})$)
23:     end for
24:  end for
In this section, we extend the result of Theorem 4.2 by taking into account the communication time $\tau_{\textnormal{w}}.$ Instead of Protocol 1, we consider Protocol 2. However, in this section, we ignore the communication times from the server to the workers and assume that $\tau_{\textnormal{s}}=0,$ which will be sufficient to obtain an almost tight lower bound if combined with Theorem 4.2.
Tyurin et al. (2024) consider a similar setup with $\tau_{\textnormal{s}}=0.$ However, their protocol does not allow the workers to modify the iterate computed by the server and operates in the primal space. For instance, the workers are not allowed to run local steps. Moreover, $\bar{P}^{k}_{i}$ are fixed in their version of Protocol 2. We improve upon this in the following theorem:
Theorem D.1.
Let $L,\Delta,\varepsilon,\sigma^{2},n,d,\tau_{\textnormal{w}},h>0$ be any numbers such that $\varepsilon<c_{1}L\Delta$ and $d\geq\frac{\Delta L}{c_{2}\varepsilon}.$ Consider Protocol 2. For all $i\in[n]$ and $k\geq 0,$ compressor $\bar{\mathcal{C}}^{k}_{i}$ selects and transmits $\bar{P}^{k}_{i}$ uniformly random coordinates without replacement, scaled by any constants, where $\bar{P}^{k}_{i}\in\{0,\dots,d\}$ may vary across each compressor. For any algorithm $A\in\mathcal{A}_{\textnormal{zr}},$ there exists a function $f\,:\,\mathbb{R}^{d}\to\mathbb{R}$ such that $f$ is $L$-smooth, $f(0)-\inf_{x\in\mathbb{R}^{d}}f(x)\leq\Delta,$ exists a stochastic gradient oracle that satisfies Assumption 1.3, and ${\mathbb{E}}\left[\inf_{y\in S_{t}}\left\|\nabla f(y)\right\|^{2}\right]>\varepsilon$ for all
@@ Appendix D Main Theorem with Worker-to-Server Communication
The proof closely follows the analysis from (Tyurin et al., 2024; Tyurin & Richtárik, 2024) and the proof of Theorem 4.2, but with some important modifications. In this proof, it is sufficient to work with (6) and Lemmas 2.1 and 2.2.
@@
where the function $F_{T}$ is given in (6) and $x_{[T]}\in\mathbb{R}^{T}$ is the vector with the first $T$ coordinates of $x\in\mathbb{R}^{d}.$ Notice that the last $d-T$ coordinates are artificial.
@@
and let $\mathcal{D}_{\xi}=\textnormal{\emph{Bernoulli}}(p_{\sigma})$ for all $j\in[n],$ where $p_{\sigma}\in(0,1].$ We denote $[x]_{j}$ as the $j$th coordinate of a vector $x\in\mathbb{R}^{d}.$ We choose
@@
seconds because $h\eta_{1,i}$ is the minimal time to discover the first coordinate locally, and $\min_{j\in[n],j\neq i}\left\{h\eta_{1,j}+\tau_{\textnormal{w}}\mu_{1,j}\right\}$ is the minimal time to discover the first coordinate from other workers via the server, which can transmit (for free since $\tau_{\textnormal{s}}=0$) it to worker $i.$
@@
For this random variable, we prove the lemma below (see Section D.1).
Lemma D.2.
@@
D.1 Main Concentration Lemma
See D.2
@@ D.1 Main Concentration Lemma
×$\quad\times\max_{i\in[n]}{\mathbb{E}}\left[\exp\left(-sy_{k-1,i}\right)\right]$

### v2

@@ D.3 Proof of lemmas
Hence, the worker receives a stream of coordinates $\nu_{1},\nu_{2},\dots$ such that $\mathbb{P}\left(\nu_{i}\in[K]\middle|\nu_{1},\dots,\nu_{i-1}\right)\leq\frac{K}{d-i+1}$ for all $i\leq d.$ Let $\mu_{1,i,1}$ be the number of received coordinates until the moment when the last received coordinate belongs to $[K]$ in worker $i.$ Similarly, let $\mu_{1,i,k}$ be the number of received coordinates until the moment when the last received coordinate belongs to $[K]$, after the (k1)(k-1)th time this has happened in worker $i.$ In other words, worker $i$ should receive $\mu_{1,i,1}$ coordinates to obtain a coordinate that belongs to $[K]$. To get the next coordinate that belongs to $[K]$, the worker should receive $\mu_{1,i,2}$ coordinates, and so on. Then,
@@
$t_{1}\geq\min_{i\in[n]}\left\{\min\left\{h\sum_{k=1}^{\frac{K}{2}}\eta_{1,i,k},\tau_{\textnormal{s}}\sum_{k=1}^{\frac{K}{2}}\mu_{1,i,k}\right\}\right\},$ (30)
@@
Remark E.2.
The previous derivations hold for all $\tau_{\textnormal{w}}>0.$ If we start taking the communication time $\tau_{\textnormal{w}}$ into account, then the bound on $t_{1}$ in (30) may only increase. For all $\tau_{\textnormal{w}}>0,$ worker $i$ still has to discover new non-zero coordinates either through stochastic gradient computations or by receiving coordinates from the server and it will take at least
$\min_{i\in[n]}\left\{\min\left\{h\sum_{k=1}^{\frac{K}{2}}\eta_{1,i,k},\tau_{\textnormal{s}}\sum_{k=1}^{\frac{K}{2}}\mu_{1,i,k}\right\}\right\}$
seconds to discover all of the first $K$ coordinates.
@@
More formally, $\eta_{2,i,k}$ can be defined as the number of stochastic gradient computations until the first moment when a coordinate is not zeroed out in (26), after the moment when the (k1)(k-1)th coordinate is no longer zeroed out, when $\textnormal{prog}^{K}$ of the input points to the stochastic gradients is $\geq K,$ and $\mu_{2,i,k}$ be the number of received coordinates until the moment when the last received coordinate belongs to $\{K+1,\dots,2K\}$, after the (k1)(k-1)th time this has happened, when $\textnormal{prog}^{1}$ of the input points to the compressor is $\geq K+1,$ and so on.
@@
E.1 Main Concentration Lemma
See E.3
@@ E.1 Main Concentration Lemma
where we bound $\max$ by $\sum.$ Using the temporal definitions of $\{\beta_{B,i}\},$
@@
Substituting the last inequality to (37) and repeating the steps B1B-1 more times, we get
@@
Appendix F Main Theorem with Worker-to-Server Communication
In this section, we extend the result of Theorem 4.2 by taking into account the communication time $\tau_{\textnormal{w}}.$ However, in this section, we ignore the communication times from the server to the workers in the analysis, which will be sufficient to obtain an almost tight lower bound if combined with Theorem 4.2.
Tyurin et al. (2024) consider a similar setup with $\tau_{\textnormal{s}}=0.$ However, their protocol does not allow the workers to modify the iterate computed by the server and operates in the primal space. For instance, the workers are not allowed to run local steps. Moreover, $\bar{P}^{k}_{i}$ are fixed in their version of Protocol 1. We improve upon this in the following theorem:
Theorem F.1.
Let $L,\Delta,\varepsilon,\sigma^{2},n,d,\tau_{\textnormal{w}},\tau_{\textnormal{s}},h>0$ be any numbers such that $\varepsilon<c_{1}L\Delta$ and $d\geq\frac{\Delta L}{c_{2}\varepsilon}.$ Consider Protocol 1. For all $i\in[n]$ and $k\geq 0,$ compressor $\bar{\mathcal{C}}^{k}_{i}$ selects and transmits $\bar{P}^{k}_{i}$ uniformly random coordinates without replacement, scaled by any constants, where $\bar{P}^{k}_{i}\in\{0,\dots,d\}$ may vary across each compressor. For any algorithm $A\in\mathcal{A}_{\textnormal{zr}},$ there exists a function $f\,:\,\mathbb{R}^{d}\to\mathbb{R}$ such that $f$ is $L$-smooth, $f(0)-\inf_{x\in\mathbb{R}^{d}}f(x)\leq\Delta,$ exists a stochastic gradient oracle that satisfies Assumption 1.3, and ${\mathbb{E}}\left[\inf_{y\in S_{t}}\left\|\nabla f(y)\right\|^{2}\right]>\varepsilon$ for all
@@ Appendix F Main Theorem with Worker-to-Server Communication
The proof closely follows the analysis from (Tyurin et al., 2024; Tyurin and Richtárik, 2024) and the proof of Theorem 4.2, but with some important modifications. In this proof, it is sufficient to work with (6) and Lemmas 2.1 and 2.2.
@@
where the function $F_{T}$ is given in (6) and $x_{[T]}\in\mathbb{R}^{T}$ is the vector with the first $T$ coordinates of $x\in\mathbb{R}^{d}.$ Notice that the last dTd-T coordinates are artificial.
@@
and let $\mathcal{D}_{\xi}=\textnormal{\emph{Bernoulli}}(p_{\sigma})$ for all $j\in[n],$ where $p_{\sigma}\in(0,1].$ We denote $x]_{j}$ as the $j$th coordinate of a vector $x\in\mathbb{R}^{d}.$ We choose
@@
seconds because $h\eta_{1,i}$ is the minimal time to discover the first coordinate locally, and $\min_{j\in[n],j\neq i}\left\{h\eta_{1,j}+\tau_{\textnormal{w}}\mu_{1,j}\right\}$ is the minimal time to discover the first coordinate from other workers via the server, which can transmit it to worker $i.$
Remark F.2.
The previous derivations hold for all $\tau_{\textnormal{s}}>0.$ If we start taking the communication time $\tau_{\textnormal{s}}$ into account, then $y_{1,i}$ may only increase. For all $\tau_{\textnormal{s}}>0,$ worker $i$ still requires at least $y_{1,i}$ seconds for the same reason that $h\eta_{1,i}$ is the minimal time to discover the first coordinate locally, and $\min_{j\in[n],j\neq i}\left\{h\eta_{1,j}+\tau_{\textnormal{w}}\mu_{1,j}\right\}$ is the minimal time to discover the first coordinate from other workers. If we start taking into account the communication time from $\tau_{\textnormal{s}},$ the lower bound
$\min\left\{h\eta_{1,i},\min_{j\in[n],j\neq i}\left\{h\eta_{1,j}+\tau_{\textnormal{w}}\mu_{1,j}\right\}\right\}$
still holds.
@@
For this random variable, we prove the lemma below (see Section F.1).
Lemma F.3.
@@
F.1 Main Concentration Lemma
See F.3
@@ F.1 Main Concentration Lemma
$\quad\times\max_{i\in[n]}{\mathbb{E}}\left[\exp\left(-sy_{k-1,i}\right)\right]$